# PointOBB-v3: Expanding Performance Boundaries of Single Point-Supervised Oriented Object Detection

### v1

@@ Abstract
With the growing demand for oriented object detection (OOD), recent studies on point-supervised OOD have attracted significant interest. In this paper, we propose PointOBB-v3, a stronger single point-supervised OOD framework. Compared to existing methods, it generates pseudo rotated boxes without additional priors and incorporates support for the end-to-end paradigm. PointOBB-v3 functions by integrating three unique image views: the original view, a resized view, and a rotated/flipped (rot/flp) view. Based on the views, a scale augmentation module and an angle acquisition module are constructed. In the first module, a Scale-Sensitive Consistency (SSC) loss and a Scale-Sensitive Feature Fusion (SSFF) module are introduced to improve the model’s ability to estimate object scale. To achieve precise angle predictions, the second module employs symmetry-based self-supervised learning. Additionally, we introduce an end-to-end version that eliminates the pseudo-label generation process by integrating a detector branch and introduces an Instance-Aware Weighting (IAW) strategy to focus on high-quality predictions. We conducted extensive experiments on the DIOR-R, DOTA-v1.0/v1.5/v2.0, FAIR1M, STAR, and RSAR datasets. Across all these datasets, our method achieves an average improvement in accuracy of 3.56% in comparison to previous state-of-the-art methods.
Index Terms:
@@ 1 Introduction
This paper is an extended version of our conference paper, published in CVPR 2024 [28], and the main contributions are as follows: 1) The proposed approach integrates object scale and orientation learning across three unique perspectives, guided by a progressive multi-view switching strategy. 2) An SSC loss is devised to strengthen the network’s capability in capturing object scale, and an SSFF module is designed to enhance accuracy in scale-related feature layer mapping. 3) An SSA loss is used to learn the object orientation and a scale-guided DS matching strategy is proposed to further enhance the precision of object angle prediction. 4) Based on the aforementioned principles, we implemented two frameworks: one end-to-end and the other two-stage, providing researchers and practitioners with a wider range of options for further exploration and application. 5) We have made improvements in the manuscript, including conducting experiments on additional benchmarks, providing deeper insights into performance improvements, conducting more comprehensive ablation studies, and reorganizing the contents and tables for better clarity [^1] This paper is built upon our work published in CVPR 2024 [28] with substantial extensions. Compared to the original version, we extend PointOBB in five aspects. 1) We propose an SSFF module to further enhance the model’s capability to perceive object scales (in Sec. 4.2). Experiments demonstrate that the proposed SSFF module significantly improves the performance of PointOBB, outperforming existing single point-supervised methods. For instance, PointOBB-v3 achieves a 2.20% improvement over PointOBB-v2 [29] on DIOR-R and an 8.76% improvement over PointOBB-v2 on DOTA-v1.0. 2) We also propose an end-to-end version of the framework that directly outputs prediction results without the need for generating pseudo-labels (in Sec. 4.4). Comprehensive experiments have been conducted to validate the effectiveness of this end-to-end framework. For example, it achieves a performance of 41.29% on the DOTA-v1.0 dataset, even surpassing two-stage methods like PointOBB. 3) We conducted additional experiments on a wider range of datasets, including DIOR, DOTA-v1.0, DOTA-v1.5, DOTA-v2.0, FAIR1M, STAR, and RSAR. Across all these datasets, our method achieves an average improvement in accuracy of 3.56% in comparison to previous state-of-the-art methods (in Sec. 5.3). 4) We provide more ablation studies to further understand the design principles and capability of PointOBB-v3. (in Sec. 5.4) 5) We have refined the manuscript by incorporating more intuitive figures and reorganizing the content and tables to enhance clarity and readability..
@@ 2.2 Weakly-Supervised Oriented Object Detection
Point-supervised. Point-based annotations have been widely used in various tasks such as object detection [45, 26, 46, 27, 24, 47], panoptic segmentation [22, 23], and instance segmentation [48, 21], among others [44, 49]. Due to its cost-effectiveness and efficiency, single point-supervised object detection has garnered significant attention. Click [27] was one of the earliest attempts at point-supervised object detection, proposing center-click annotations and incorporating them into a MIL paradigm. P2BNet [26] further enhances this approach by using a coarse-to-fine strategy and incorporating negative samples to improve prediction quality. However, these methods only produce horizontal bounding boxes and fail to address the inherent instability of the MIL paradigm effectively.
@@ 4.1.2 Progressive Multi-View Switching Strategy
This progressive multi-view switching strategy operates through three stages: Stage 1. The process begins by generating a resized view from the original view using a scaling factor $\sigma$. This resized view, combined with the original view, adheres to scale-equivalence constraints, forming the basis of the scale augmentation module. This module is designed to improve the network’s capability to accurately perceive object scales, laying the groundwork for effective scale-sensitive learning. Details of this module will be discussed in subsequent sections. Stage 2 (i.e., “burn-in step1” in Fig. 2). Before this stage, the network has developed fundamental perceptual abilities to the scale and boundary of objects. However, orientation information is still lack. To address this, the resized view is replaced with a rot/flp view, which is paired with the original view to construct the angle acquisition module. This module employs a dense-to-dense sample assignment approach, enabling the network to engage in self-supervised angle learning, thereby introducing orientation sensitivity. Stage 3 (i.e., “burn-in step2” in Fig. 2). At this stage, the network has achieved a refined ability to predict accurate angles using dense feature representations. To integrate these angle predictions with object proposals, the proposed DS matching strategy is employed. This strategy aligns dense angle predictions with sparse proposals by leveraging neighboring receptive fields, effectively assigning accurate object orientations to the proposals.
@@ 4.2.1 Scale-Sensitive Consistency Loss
$\text{sim}_{m,g}^{\text{ins}}=1-\frac{\left[S_{\text{i{o}}}^{\text{ins}}\right]_{m,g}\cdot\left[S_{\text{i{d}}}^{\text{ins}}\right]_{m,g}}{\left\|\left[S_{\text{i{o}}}^{\text{ins}}\right]_{m,g}\right\|\cdot\left\|\left[S_{\text{i{d}}}^{\text{ins}}\right]_{m,g}\right\|},$ (4)
$\text{sim}_{m,g}^{\text{cls}}=1-\frac{\left[S_{\text{i{o}}}^{\text{cls}}\right]_{m,g}\cdot\left[S_{\text{i{d}}}^{\text{cls}}\right]_{m,g}}{\left\|\left[S_{\text{i{o}}}^{\text{cls}}\right]_{m,g}\right\|\cdot\left\|\left[S_{\text{i{d}}}^{\text{cls}}\right]_{m,g}\right\|},$
@@
$\mathcal{L}_{\text{SSC}}=\sum_{m=1}^{M}\sum_{g=1}^{G}\left\{\omega_{1}\ell_{s}\left(\text{sim}_{m,g}^{\text{ins}},0\right)+\omega_{2}\ell_{s}\left(\text{sim}_{m,g}^{\text{cls}},0\right)\right\},$ (5)
@@ 4.2.2 Scale-Sensitive Feature Fusion Module
Such a design can automatically deal with ROIs of different sizes. Most importantly, this module ensures that the final extracted ROI features contain more accurate scale information. Adopting this more flexible method helps mitigate issues such as mapping deviation and inconsistencies in feature extraction caused by inaccuracies in proposal scales. Taking advantage of this novel mechanism SSFF, PointOBB-v3 achieves a much higher AP50 compared to our conference work PointOBB [28] (40.18% vs 37.31% on DIOR-R testing set and 50.44% vs 33.31% on DOTA-v1.0 testing set, see in Tab. IV).
@@ 4.3.3 Self-Supervised Angle Loss
Leveraging the affine transformation relationship between the original view and the rotated/flipped view, object angles are learned using the Self-Supervised Angle (SSA) loss. If the enhanced view is created through a random rotation by an angle $\theta^{\prime}$, the angle predictions from both the original and rotated views should align with and uphold the same rotational relationship. When the enhanced view is produced through a vertical flip, the angle predictions must account for differences of $k\pi$, where $k$ is an integer ensuring the results remain within the same periodic cycle. The loss between the outputs of two views can be formulated as:
@@ 4.4.1 Pipeline
With the above pipeline design, the model can be trained and inferred in a more convenient and time-efficient end-to-end manner. As demonstrated in Tab. III , compared to the two-stage version, the end-to-end version reduces the training hours by approximately 21.36%, while maintaining competitive detection performance.
TABLE I: Accuracy on the DIOR-R testing set.
@@
Comparison tracks: ✓= End-to-end training and testing; $\times$ = Generating pseudo labels to train the detector (two-stage training).
The term “e2e” refers to the end-to-end version of PointOBB-v3.
APL: Airplane, APO: Airport, BF: Baseball Field, BC: Basketball Court, BR: Bridge, CH: Chimney, ESA: Expressway Service Area,
ETS: Expressway Toll Station, DAM: Dam, GF: Golf Field, GTF: Ground Track Field, HA: Harbor, OP: Overpass, SH: Ship,
STA: Stadium, STP: Storage Tank, TC: Tennis Court, TS: Train Station, VE: Vehicle, WM: Windmill
-RBOX: The minimum rectangle operation is performed on the Mask to obtain RBox.
4.4.2 Instance-Aware Weighting Strategy
4.5 Optimization with Overall MIL Loss
In detail, each point label corresponds to a specific proposal bag. Using the MIL head, the instance score and class score for the $n$-th proposal within the $i$-th proposal bag are represented as $S_{\textit{i,n}}^{ins}$ and $S_{\textit{i,n}}^{cls}$, respectively. The overall output for the entire proposal bag is expressed as $S_{i}=\sum_{n=1}^{N}S_{\textit{i,n}}^{ins}\odot S_{\text{i,n}}^{cls}$, where $N$ indicates the total number of proposals in the bag. The MIL losses corresponding to the original view, rotated/flipped view, and resized view are represented as $\mathcal{L}_{\text{MIL}}^{\text{ori}}$, $\mathcal{L}_{\text{MIL}}^{\text{rfv}}$, and $\mathcal{L}_{\text{MIL}}^{\text{res}}$, respectively. These three losses follow the common form of the general MIL loss $\mathcal{L}_{\text{MIL}}^{\text{init}}$, which is defined as:
$\mathcal{L}_{\text{MIL}}^{\text{init}}=$ $-\frac{1}{I}\sum_{i=1}^{I}\sum_{c=1}^{C}\{Q_{i,c}\log(S_{i,c})$ (11)
$+(1-Q_{i,c})(1-\log(1-S_{i,c}))\},$
where $\alpha$ and $\beta$ keep the same setting with Eq. 13.
5 Experiments
5.1 Datasets and Implementation Details
TABLE II: Accuracy on the DOTA-v1.0 testing set.
@@ 5.1 Datasets and Implementation Details
PointOBB (FCOSR) [28] $\times$ 26.1 65.7 9.1 59.4 65.8 34.9 29.8 0.5 2.3 16.7 0.6 49.0 21.8 41.0 36.7 30.08
@@
Comparison tracks: ✓= End-to-end training and testing; $\times$ = Generating pseudo labels to train the detector (two-stage training).
The term “e2e” refers to the end-to-end version of PointOBB-v3.
PL: Plane, BD: Baseball diamond, BR: Bridge, GTF: Ground track field, SV: Small vehicle, LV: Large vehicle, SH: Ship, TC: Tennis court,
BC: Basketball court, ST: Storage tank, SBF: Soccer-ball field, RA: Roundabout, HA: Harbor, SP: Swimming pool, HC: Helicopter.
TABLE III: Comparison of the training time and the accuracy between the two-stage version and the end-to-end version on the DIOR-R dataset.
Method Epochs Train hours mAP50
Two-stage 36 17.23 40.18
End-to-end 24 13.55 37.60
TABLE IV: The effects of SSC loss, DS matching strategy and SSFF in DIOR-R and DOTA-v1.0 datasets. We evaluate Rotated FCOS on the DIOR-R testing set and Oriented R-CNN on the DOTA-v1.0 testing set.
Module DIOR-R [65] DOTA-v1.0 [25]
SSC DS SSFF mIoU mAP50 mIoU mAP50
47.95 30.16 40.26 28.44
50.78 31.96 42.54 30.63
53.17 36.39 43.92 32.98
56.08 37.31 45.35 33.31
57.68 40.18 51.33 50.44
Experiments are performed on a server with 2 A100 GPUs and 80GB memory. For the two-stage version, we use the “2×” schedule to train all methods and apply the “1×” schedule when training RBox-supervised algorithms with our generated pseudo OBBs. For the end-to-end version, the entire detector is trained using the “2×” schedule. This paper inherits PointOBB’s optimal default settings, employing the SGD optimizer with a learning rate of 0.005, momentum of 0.9, and weight decay of 0.0001. A linear warm-up strategy is applied for the initial 500 iterations with a rate of 0.001 and batch size is 2. With the exception of WSODet [41], which utilizes VGG16 [72], all the listed models are built on the ResNet50 [63] backbone, pre-trained on ImageNet [73]. For the comparative algorithms, their default settings are adopted. In our method, for both the two-stage version and the end-to-end version, the finest scale is set to 56, and the scale factor $\sigma$ for the resized view is randomly selected between 0.5 and 1.5. Following the settings in PointOBB, “burn-in step1” and “burn-in step2” are configured at the 6th and 8th epochs, respectively. Additionally, we adopt scale-based grouping, which has been demonstrated through experiments in PointOBB to be more effective than proposal-based and ratio-based grouping strategies.
@@ 5.2.1 Results on DIOR-R
Tab. I compares PointOBB-v3 with existing methods, which can be categorized into two tracks:
TABLE V: Results on more datasets.
@@
Results on DIOR, DOTA-v1.0/v1.5/v2.0, FAIR1M, STAR and RSAR datasets, reporting the mAP50 metric.
The term “e2e” refers to the end-to-end version of PointOBB-v3.
* indicates using additional human knowledge priors.
2) End-to-end mode. This paradigm directly trains a weakly supervised detector on the training and validation sets and performs inference on the test set. Without utilizing any priors, our approach achieves a 10.26% improvement (37.60% vs. 27.34%) over Point2RBox [50]. Remarkably, it also demonstrates competitive performance compared to methods in the two-stage mode. For instance, it surpasses PointOBB (FCOSR) by 0.29% and PointOBB-v2 [29] (FCOSR) by 3.25%, respectively. Furthermore, as shown in Tab. III, while maintaining competitive performance, the end-to-end version also achieves improvements in efficiency. Specifically, the number of training epochs is reduced from 36 to 24, and the total training time decreases from 17.23 hours to 13.55 hours—a reduction of approximately 21.36%.
5.2.2 Results on DOTA-v1.0
As shown in Tab. II, the comparison of PointOBB-v3 with existing methods is organized into two distinct tracks:
2) End-to-end mode. The end-to-end version of PointOBB-v3 attains AP50 scores of 41.29% on DOTA-v1.0. Notably, it outperforms the two-stage counterparts, PointOBB and PointOBB-v2. Specifically, it achieves an improvement of 11.21% over PointOBB (FCOSR) and maintains comparable performance to PointOBB-v2 across its two detectors. Additionally, when compared to Point2RBox, which leverages manual sketches to assist in boundary determination, our approach achieves a remarkable gain of 1.02%.
5.3 Results on More Datasets
The results are displayed in Tab. V. On the more challenging DOTA-v1.5 dataset, PointOBB-v3 exhibits a comparable trend, outperforming PointOBB by 23.13%/27.16% and surpassing PointOBB-v2 by 3.2%/6.07% in the pseudo-generation track. Similarly, on DOTA-v2.0, PointOBB-v3 achieves a 17.99%/18.57% improvement over PointOBB and a 2.88%/1.46% gain over PointOBB-v2. Our method also demonstrates strong performance on various datasets such as FAIR1M, STAR and RSAR. Specifically, PointOBB-v3 achieves improvements of 7.16%/3.66%/8.80% compared to PointOBB and 4.99%/3.85%/3.61% over PointOBB-v2 with the Rotated FCOS detector.
On the other hand, the end-to-end version of PointOBB-v3 also delivers strong performance on these more challenging datasets. Specifically, it outperforms Point2RBox-RC on DOTA-v1.5, DOTA-v2.0, FAIR1M, and STAR datasets. Even when compared to Point2RBox-SK, which incorporates human prior knowledge, our method also demonstrates highly competitive performance, particularly achieving a certain degree of superiority on the DIOR, DOTA-v1.0, and DOTA-v1.5 datasets (37.60% vs. 27.34%, 41.29% vs. 40.27%, 31.25% vs. 30.51%). Moreover, on the STAR dataset, our end-to-end version outperforms PointOBB (FCOSR) and PointOBB-v2 (FCOSR) by 2.12% and 2.31%, respectively, while achieving gains of 2.04% and 4.28% on the RSAR dataset.
5.4 Ablation Studies
TABLE VI: Ablation studies of using MIL loss at different views on DIOR-R. mAP is reported by training the Oriented R-CNN. “O” indicates the original view, “R” indicates the resized view, “R/F” indicates the Rot/Flp view.
@@ 5.4 Ablation Studies
TABLE VII: Ablation studies of the burn-in steps on DIOR-R. mAP is reported by training the Oriented R-CNN. “P”: Plainly, “G”: Gradually, “A” indicates attaching gradients between the angle branch and the other parts.

### v2

# latexfontFont

@@ Abstract
With the growing demand for oriented object detection (OOD), recent studies on point-supervised OOD have attracted significant interest. In this paper, we propose PointOBB-v3, a stronger single point-supervised OOD framework. Compared to existing methods, it generates pseudo rotated boxes without additional priors and incorporates support for the end-to-end paradigm. PointOBB-v3 functions by integrating three unique image views: the original view, a resized view, and a rotated/flipped (rot/flp) view. Based on the views, a scale augmentation module and an angle acquisition module are constructed. In the first module, a Scale-Sensitive Consistency (SSC) loss and a Scale-Sensitive Feature Fusion (SSFF) module are introduced to improve the model’s ability to estimate object scale. To achieve precise angle predictions, the second module employs symmetry-based self-supervised learning. Additionally, we introduce an end-to-end version that eliminates the pseudo-label generation process by integrating a detector branch and introduces an Instance-Aware Weighting (IAW) strategy to focus on high-quality predictions. We conducted extensive experiments on the DIOR-R, DOTA-v1.0/v1.5/v2.0, FAIR1M, STAR, and RSAR datasets. Across all these datasets, our method achieves an average improvement in accuracy of 3.56% in comparison to previous state-of-the-art methods. The code will be available at https://github.com/ZpyWHU/PointOBB-v3.
keywords:
@@ 1 Introduction
This paper is an extended version of our conference paper, published in CVPR 2024 [28], and the main contributions are as follows  [^1] This paper is built upon our work published in CVPR 2024 [28] with substantial extensions. Compared to the original version, we extend PointOBB in five aspects. 1) We propose an SSFF module to further enhance the model’s capability to perceive object scales (in Sec. 4.2). Experiments demonstrate that the proposed SSFF module significantly improves the performance of PointOBB, outperforming existing single point-supervised methods. For instance, PointOBB-v3 achieves a 2.20% improvement over PointOBB-v2 [29] on DIOR-R and an 8.76% improvement over PointOBB-v2 on DOTA-v1.0. 2) We also propose an end-to-end version of the framework that directly outputs prediction results without the need for generating pseudo-labels (in Sec. 4.4). Comprehensive experiments have been conducted to validate the effectiveness of this end-to-end framework. For example, it achieves a performance of 41.29% on the DOTA-v1.0 dataset, even surpassing two-stage methods like PointOBB. 3) We conducted additional experiments on a wider range of datasets, including DIOR, DOTA-v1.0, DOTA-v1.5, DOTA-v2.0, FAIR1M, STAR, and RSAR. Across all these datasets, our method achieves an average improvement in accuracy of 3.56% in comparison to previous state-of-the-art methods (in Sec. 5.3). 4) We provide more ablation studies to further understand the design principles and capability of PointOBB-v3. (in Sec. 5.4) 5) We have refined the manuscript by incorporating more intuitive figures and reorganizing the content and tables to enhance clarity and readability.: 1) The proposed approach integrates object scale and orientation learning across three unique perspectives, guided by a progressive multi-view switching strategy. 2) An SSC loss is devised to strengthen the network’s capability in capturing object scale, and an SSFF module is designed to enhance accuracy in scale-related feature layer mapping. 3) An SSA loss is used to learn the object orientation and a scale-guided DS matching strategy is proposed to further enhance the precision of object angle prediction. 4) Based on the aforementioned principles, we implemented two frameworks: one end-to-end and the other two-stage, providing researchers and practitioners with a wider range of options for further exploration and application. 5) We have made improvements in the manuscript, including conducting experiments on additional benchmarks, providing deeper insights into performance improvements, conducting more comprehensive ablation studies, and reorganizing the contents and tables for better clarity.
@@ 2.2 Weakly-Supervised Oriented Object Detection
Point-supervised. Point-based annotations have been widely used in various tasks such as object detection [45, 26, 46, 27, 24, 47, 48], panoptic segmentation [22, 23], and instance segmentation [49, 21], among others [44, 50]. Due to its cost-effectiveness and efficiency, single point-supervised object detection has garnered significant attention. Click [27] was one of the earliest attempts at point-supervised object detection, proposing center-click annotations and incorporating them into an MIL paradigm. P2BNet [26] further enhances this approach by using a coarse-to-fine strategy and incorporating negative samples to improve prediction quality. However, these methods only produce horizontal bounding boxes and fail to address the inherent instability of the MIL paradigm effectively.
@@ 4.1.2 Progressive Multi-View Switching Strategy
This progressive multi-view switching strategy operates through three stages: Stage 1. The process begins by generating a resized view from the original view using a scaling factor $\sigma$. This resized view, combined with the original view, adheres to scale-equivalence constraints, forming the basis of the scale augmentation module. This module is designed to improve the network’s capability to accurately perceive object scales, laying the groundwork for effective scale-sensitive learning. Details of this module will be discussed in subsequent sections. Stage 2 (i.e., “burn-in step1” in Fig. 2). Before this stage, the network has developed fundamental perceptual abilities to the scale and boundary of objects. However, orientation information is still lacking. To address this, the resized view is replaced with a rot/flp view, which is paired with the original view to construct the angle acquisition module. This module employs a dense-to-dense sample assignment approach, enabling the network to engage in self-supervised angle learning, thereby introducing orientation sensitivity. Stage 3 (i.e., “burn-in step2” in Fig. 2). At this stage, the network has achieved a refined ability to predict accurate angles using dense feature representations. To integrate these angle predictions with object proposals, the proposed DS matching strategy is employed. This strategy aligns dense angle predictions with sparse proposals by leveraging neighboring receptive fields, effectively assigning accurate object orientations to the proposals.
@@ 4.2.1 Scale-Sensitive Consistency Loss
[…]
@@
$\mathcal{L}_{\textit{SSC}}=\sum_{m=1}^{M}\sum_{g=1}^{G}\left\{\omega_{1}\ell_{s}\left(\textit{sim}_{m,g}^{\textit{ins}},0\right)+\omega_{2}\ell_{s}\left(\textit{sim}_{m,g}^{\textit{cls}},0\right)\right\},$ (6)
@@ 4.2.2 Scale-Sensitive Feature Fusion Module
Such a design can automatically deal with ROIs of different sizes. Most importantly, this module ensures that the final extracted ROI features contain more accurate scale information. Adopting this more flexible method helps mitigate issues such as mapping deviation and inconsistencies in feature extraction caused by inaccuracies in proposal scales. Taking advantage of this novel mechanism SSFF, PointOBB-v3 achieves a much higher AP50 compared to our conference work PointOBB [28] (40.18% vs 37.31% on DIOR-R testing set and 50.44% vs 33.31% on DOTA-v1.0 testing set, see in Tab. 5).
@@ 4.3.3 Self-Supervised Angle Loss
Leveraging the affine transformation relationship between the original view and the rotated/flipped view, object angles are learned using the Self-Supervised Angle (SSA) loss. If the enhanced view is created through a random rotation by an angle $\theta^{\prime}$, the angle predictions from both the original and rotated views should align with and uphold the same rotational relationship. When the enhanced view is produced through a vertical flip, the angle predictions must account for differences of $k\pi$, where $k$ is an integer ensuring the results remain within the same periodic cycle. The loss between the outputs of the two views can be formulated as:
@@ 4.4.1 Pipeline
With the above pipeline design, the model can be trained and inferred in a more convenient and time-efficient end-to-end manner. As demonstrated in Tab. 3 , compared to the two-stage version, the end-to-end version reduces the training hours by approximately 21.36%, while maintaining competitive detection performance.
4.4.2 Instance-Aware Weighting Strategy
4.5 Optimization with Overall MIL Loss
In detail, each point label corresponds to a specific proposal bag. Using the MIL head, the instance score and class score for the $n$-th proposal within the $i$-th proposal bag are represented as $S_{\textit{i,n}}^{ins}$ and $S_{\textit{i,n}}^{cls}$, respectively. The overall output for the entire proposal bag is expressed as $S_{i}=\sum_{n=1}^{N}S_{\textit{i,n}}^{ins}\odot S_{\textit{i,n}}^{cls}$, where $N$ indicates the total number of proposals in the bag. The MIL losses corresponding to the original view, rotated/flipped view, and resized view are represented as $\mathcal{L}_{\text{MIL}}^{\text{ori}}$, $\mathcal{L}_{\text{MIL}}^{\text{rfv}}$, and $\mathcal{L}_{\text{MIL}}^{\text{res}}$, respectively. These three losses follow the common form of the general MIL loss $\mathcal{L}_{\text{MIL}}^{\text{init}}$, which is defined as:
$\mathcal{L}_{\text{MIL}}^{\text{init}}=-\frac{1}{I}\sum_{i=1}^{I}\sum_{c=1}^{C}\left\{Q_{i,c}\log(S_{i,c})+(1-Q_{i,c})(1-\log(1-S_{i,c}))\right\},$ (12)
where $\alpha$ and $\beta$ keep the same setting with Eq. 14.
5 Experiments
5.1 Datasets and Implementation Details
Experiments are performed on a server with 2 A100 GPUs and 80GB memory. For the two-stage version, we use the “2×” schedule to train all methods and apply the “1×” schedule when training RBox-supervised algorithms with our generated pseudo OBBs. For the end-to-end version, the entire detector is trained using the “2×” schedule. This paper inherits PointOBB’s optimal default settings, employing the SGD optimizer with a learning rate of 0.005, momentum of 0.9, and weight decay of 0.0001. A linear warm-up strategy is applied for the initial 500 iterations with a rate of 0.001 and batch size is 2. With the exception of WSODet [41], which utilizes VGG16 [73], all the listed models are built on the ResNet50 [64] backbone, pre-trained on ImageNet [74]. For the comparative algorithms, their default settings are adopted. In our method, for both the two-stage version and the end-to-end version, the finest scale is set to 56, and the scale factor $\sigma$ for the resized view is randomly selected between 0.5 and 1.5. Following the results of the ablation study, “burn-in step1” and “burn-in step2” are configured at the 6th and 8th epochs, respectively. Additionally, we adopt scale-based grouping, which has been demonstrated through experiments in PointOBB to be more effective than proposal-based and ratio-based grouping strategies.
Table 1: Accuracy on the DIOR-R testing set.
@@ 5.1 Datasets and Implementation Details
[…]
Table 2: Accuracy on the DOTA-v1.0 testing set.
@@
PointOBB (FCOSR) [28] $\times$ 26.1 65.7 9.1 59.4 65.8 34.9 29.8 0.5 2.3 16.7 0.6 49.04 21.8 41.0 36.7 30.08
@@
[…]
@@ 5.2.1 Results on DIOR-R
Tab. 2 compares PointOBB-v3 with existing methods, which can be categorized into two tracks:
2) End-to-end mode. This paradigm directly trains a weakly supervised detector on the training and validation sets and performs inference on the test set. Without utilizing any priors, our approach achieves a 10.26% improvement (37.60% vs. 27.34%) over Point2RBox [51]. Remarkably, it also demonstrates competitive performance compared to methods in the two-stage mode. For instance, it surpasses PointOBB (FCOSR) by 0.29% and PointOBB-v2 [29] (FCOSR) by 3.25%, respectively. Furthermore, as shown in Tab. 3, while maintaining competitive performance, the end-to-end version also achieves improvements in efficiency. Specifically, the number of training epochs is reduced from 36 to 24, and the total training time decreases from 17.23 hours to 13.55 hours—a reduction of approximately 21.36%.
3) Computational efficiency. As shown in Tab. 11, although PointOBB-v3 does have a longer training time compared to other methods, the significant performance improvement compensates for the additional training cost. Since PointOBB-v3 is an extension of PointOBB, it maintains a similar level of efficiency while achieving notable performance improvements. In contrast, extending the training time to 36 epochs on Point2RBox-SK and PointOBB-v2 yields almost no significant performance improvement. Additionally, in terms of inference speed (pseudo-label generation), PointOBB-v3 significantly outperforms PointOBB-v2, especially since PointOBB-v2 requires PCA computation during inference to generate pseudo-labels, which significantly slows down its overall inference speed.
Table 3: Comparison of the training time and the accuracy between the two-stage version and the end-to-end version on the DIOR-R dataset.
Method Epochs Train hours mAP50
Two-stage 36 17.23 40.18111The categories in DIOR-R include Airplane (APL), Airport (APO), Baseball Field (BF), Basketball Court (BC), Bridge (BR), Chimney (CH), Expressway Service Area (ESA), Expressway Toll Station (ETS), Dam (DAM), Golf Field (GF), Ground Track Field (GTF), Harbor (HA), Overpass (OP), Ship (SH), Stadium (STA), Storage Tank (STO), Tennis Court (TC), Train Station (TS), Vehicle (VE) and Windmill (WM). ”-RBox” means the minimum rectangle operation is performed on the Mask to obtain RBox. * Comparison tracks: ✓= End-to-end training and testing; $\times$ = Generating pseudo labels to train the detector (two-stage training).The categories in the DOTA-v1.0 dataset include Plane (PL), Baseball Diamond (BD), Bridge (BR), Ground Track Field (GTF), Small Vehicle (SV), Large Vehicle (LV), Ship (SH), Tennis Court (TC), Basketball Court (BC), Storage Tank (ST), Soccer Ball Field (SBF), Roundabout (RA), Harbor (HA), Swimming Pool (SP), and Helicopter (HC). * Comparison tracks: ✓= End-to-end training and testing; $\times$ = Generating pseudo labels to train the detector (two-stage training).
End-to-End 24 13.55 37.60
[…]
11footnotetext: The reported mAP50 is trained with Rotated FCOS.
(a) SSA loss
(b) MIL loss
Figure 6: (a) indicates the SSA loss under different settings. In (b), “MIL loss (Gradually)” means training the network with the proposed gradual strategy (i.e., in chronological order), and “MIL loss (Plainly)” means training the angle branch from the beginning directly.
5.2.2 Results on DOTA-v1.0
As shown in Tab. 2, the comparison of PointOBB-v3 with existing methods is organized into two distinct tracks:
2) End-to-end mode. The end-to-end version of PointOBB-v3 attains AP50 scores of 41.29% on DOTA-v1.0. Specifically, it achieves an improvement of 11.21% over PointOBB (FCOSR) and maintains comparable performance to PointOBB-v2 across its two detectors. Additionally, when compared to Point2RBox, which leverages manual sketches to assist in boundary determination, our approach achieves a remarkable gain of 1.02%.
Table 4: Results on more datasets.
@@
11footnotetext: Results on DIOR, DOTA-v1.0/v1.5/v2.0, FAIR1M, STAR and RSAR datasets, reporting the mAP50 metric. The term e2e refers to the end-to-end version of the framework. Rotated FCOS and Oriented R-CNN refer to the two-stage version of the framework, using Rotated FCOS and Oriented R-CNN as detectors, respectively. * indicates using additional human knowledge priors.
Table 5: The effects of SSC loss, DS matching strategy and SSFF in DIOR-R and DOTA-v1.0 datasets. We evaluate Rotated FCOS on the DIOR-R testing set and Oriented R-CNN on the DOTA-v1.0 testing set.
Module DIOR-R [66] DOTA-v1.0 [25]
SSC DS SSFF mIoU mAP50 mIoU mAP50
47.95 30.16 40.26 28.44
50.78 31.96 42.54 30.63
53.17 36.39 43.92 32.98
56.08 37.31 45.35 33.31
57.68 40.18 51.33 50.44
Table 6: Ablation studies of using MIL loss at different views on DIOR-R. mAP is reported by training the Oriented R-CNN. “O” indicates the original view, “R” indicates the resized view, “R/F” indicates the Rot/Flp view.
@@
Table 7: Ablation studies of the burn-in steps on DIOR-R. mAP is reported by training the Oriented R-CNN. “P”: Plainly, “G”: Gradually, “A” indicates attaching gradients between the angle branch and the other parts.